# Neighbour-Driven Gaussian Process Variational Autoencoders for Scalable Structured Latent Modelling

### v2

@@ 1 Introduction
[…]
@@ 3.2 Neighbour-Driven Inference
where $\mathbf{D}_{\mathbf{w}}=\operatorname{diag}(\mathbf{w})$. From (6), the correlations of the unselected variables $\absent\mathbf{Z}_{\backslash\mathcal{I}}$ among $\mathbf{Z}$ are eliminated, where \=$\backslash\mathcal{I}=\{i:\mathbf{w}_{i}=0\}$. Then, we set up an amortised variational distribution that uses a similar form to (6) as follows [^1] We have omitted the conditioning variable $\mathbf{Y}$.
@@
.$\quad\quad\quad\quad\biggl{.}-\frac{1}{N}\mathrm{KL}\left[q(\mathbf{Z}_{n(i)})\mid\mid p(\mathbf{Z}_{n(i)})\right]\biggr{\}},$
@@
[…]
@@ 3.3 Predictive Posterior
Predictive posterior distributions $p(\mathbf{y}_{*}\mid\mathbf{x}_{*},\mathbf{Y})$ are often used in tasks of conditional generation, where an unseen auxiliary $\mathbf{x}_{*}\in\mathbb{R}^{D}$ is given. Here, we derive an approximate predictive posterior for the proposed GPVAEs based on neighbouring latent representations. Specifically, computing $p(\mathbf{y}_{*}\mid\mathbf{x}_{*},\mathbf{Y})$ at a new location $\mathbf{x}_{*}$ only needs to consider its $H$ nearest neighbours in $\mathbf{X}$ (denoted by the index set $n(*)$):
@@ 5.2 Rotated MNIST
Corrupted Frames VAE (Kingma & Welling, 2014) 0.193 $\pm$ 0.038 0.333 $\pm$ 0.011 75.741 $\pm$ 0.660
@@
Missing Frames SVGPVAE-M10 (Jazbec et al., 2021) 87.711 $\pm$ 0.484 6.122 $\pm$ 0.015 8.280 $\pm$ 0.111
@@
[…]
@@ 5.4 Geostatistical datasets
Imputation results are summarised in Table 3. We present plots that illustrate the imputation of porosity at the intermediate layer. In this highly heterogeneous dataset, methods that rely on inducing points, such as Multi-Output GPs (MOGP), SVGPVAE, and SGPBAE, all perform poorly, even using a large number of inducing points (e.g., MOGP-M1000). This is because inducing point methods struggle to effectively manage data which exhibit rapid local changes. In contrast, models employing nearest neighbour approximation (including our models and VNNGP) generally perform better on this dataset, indicating that the nearest neighbour approach helps capture local structure. Importantly, our models outperform VNNGP with  15$\times$ fewer trainable parameters, demonstrating that the VAE structure effectively models the latent correlations in the data.
@@ Acknowledgements
We thank the reviewers for their insightful comments and constructive suggestions. We also acknowledge the CSF3 at the University of Manchester for providing GPU resources. Xinxing is supported by the UoM-CSC Joint Scholarship, and Xiaoyu receives support through UoM Departmental Studentship for Department of Computer Science.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ B.1 GPVAE-HPA ELBO
We begin by outlining the derivation for GPVAE-HPA. In each training iteration, randomly sample a mini-batch =$\mathcal{B}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i\in\mathcal{I}}$ with the index set $\mathcal{I}$. For each training point $(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathcal{B}$, look for the $H$ locations $\mathbf{X}_{n(i)}\in\mathbb{R}^{H\times D}$ in $\mathbf{X}=[\mathbf{x}_{n}]_{n=1}^{N}$ such that $n(i)$ are indices of top $H$ nearest points to $\mathbf{x}_{i}$. This neighbour-driven strategy materialises the sampling of a binary indicative vector $\mathbf{w}$ for latent variables. Then, the ELBO is given by
@@ B.2 GPVAE-SPA ELBO
    $\mathbf{Z}_{n(j)}\sim\prod_{l=1}^{L}q_{\phi}(\mathbf{Z}^{l}_{n(j)}\mid\mathbf{Y}_{n(j)})=\prod_{l=1}^{L}\mathcal{N}\left(\mathbf{Z}^{l}_{n(j)}\mid\boldsymbol{\mu}^{l}_{n(j)},\mathbf{S}^{l}_{n(j)}\right),$
@@ C.1 Moving Ball
Frame size 32$\times$ 32
@@ C.2.2 Missing Frame Generation
Decoder 16 $\rightarrow$ 1568 $\rightarrow$ Deconv(32, 64, ks=3) $\rightarrow$ Deconv(64, 32, ks=3) $\rightarrow$ Deconv(32, 1, ks=3) $\rightarrow$ Sigmoid
@@ C.4.1 Jura
[…]
@@ Appendix D Scalable Gaussian Process Variational Autoencoders
Table 15: A summary of relevant models. $N$, $M$, and $H$ are the number of data points, inducing points, and nearest neighbours. $N_{b}$ is the mini-batch size. $P$ represents the feature dimensions of low-rank representations. $d$ is the dimensions of the state variables.
Models Mini-batching* GP complexity Arbitrary kernel Standard VAE encoder Reference
@@ D.1 SVGPVAE
Replacing exact GPs with sparse variational GP (SVGP) techniques (Titsias, 2009; Hensman et al., 2013) to enhance the scalability of GPVAEs is a natural approach. However, such methods cannot be directly integrated into GPVAEs, as they either require the entire dataset to be loaded into memory or are not suited for amortization. In response to this scalability challenge, Jazbec et al. (2021) propose a novel approach that leverages a specific parametrisation of the distribution of inducing variables based on Pearce (2020). This formulation enables the training objective to accommodate mini-batching, amortization, and the use of arbitrary kernels.
@@
where $Z^{l}_{\psi,\phi}=\mathcal{N}(\mu^{l}_{\phi}(\mathbf{Y})\mid\mathbf{0},k^{l}_{\psi}(\mathbf{X},\mathbf{X})+{\sigma^{2}_{\phi}}^{l}(\mathbf{Y}))$. The posterior $q_{\psi,\phi}(\mathbf{z}_{n}\mid\mathbf{y}_{n})$ can be regarded as derived from a “pseudo” GP regerssion with input-observation pairs $(\mathbf{X},\mu^{l}_{\phi}(\mathbf{Y}))$ and noise ${\sigma_{\phi}^{2}}^{l}(\mathbf{Y})$. We denote the latent dataset as $\{\mathbf{X},\tilde{\mathbf{Y}}=\mu_{\phi}(\mathbf{Y}),\tilde{\boldsymbol{\sigma}}^{2}={\sigma_{\phi}^{2}}^{l}(\mathbf{Y})\}$. This training objective will require $\mathcal{O}(N^{3})$ time complexity due to $\log Z^{l}_{\psi,\phi}(\mathbf{Y},\mathbf{X})$ and $q_{\psi,\phi}(\mathbf{z}_{n}\mid\mathbf{y}_{n})$ terms without any approximation. Jazbec et al. (2021) approximate the encoder’s output distribution $q_{\psi,\phi}(\mathbf{z})$ using SVGP’s posterior:
@@
$\mathbf{\Sigma}_{b}^{l}$ $:=\mathbf{K_{UU}}+\frac{N}{N_{b}}\mathbf{K}_{\mathbf{UX}_{b}}\operatorname{diag}(\tilde{\boldsymbol{\sigma}}_{b}^{-2})\mathbf{K}_{\mathbf{X}_{b}\mathbf{U}},$ (18)
$\boldsymbol{\mu}_{b}^{l}$ $:=\frac{N}{N_{b}}\mathbf{K_{UU}}(\mathbf{\Sigma}_{b}^{l})^{-1}\mathbf{K}_{\mathbf{U}\mathbf{X}_{b}}\operatorname{diag}(\tilde{\boldsymbol{\sigma}}_{b}^{-2})\tilde{\mathbf{Y}}^{l}_{b},$
$\mathbf{A}_{b}^{l}$ $:=\mathbf{K_{UU}}(\mathbf{\Sigma}_{b}^{l})^{-1}\mathbf{K_{UU}}.$
@@
[…]
The terms in the above objective are specified as
$q_{s}(\mathbf{z}_{n})=\prod_{l=1}^{L}q_{s}(\mathbf{z}^{l}_{n}),$
[…]
$\mathcal{L}^{l}_{H}&=\sum_{n=1}^{N}\left\{\log\mathcal{N}\left(\tilde{y}^{l}_{n}\mid\mathbf{b}^{l}_{b}(\mathbf{x}_{n})\boldsymbol{\mu}_{b}^{l},\tilde{\sigma}^{2}_{n}\right)-\frac{1}{2\tilde{\sigma}^{2}_{n}}\left[\tilde{k}_{nn}+\operatorname{tr}(\mathbf{A}^{l}_{b}\boldsymbol{\Lambda}_{n})\right]\right\}-\text{KL}[q(\mathbf{z}^{l}_{\mathbf{U}})\mid\mid p_{\psi}(\mathbf{z}^{l}_{\mathbf{U}})]\\ &\approx\frac{N}{N_{b}}\sum_{n=1}^{N_{b}}\left\{\log\mathcal{N}\left(\tilde{y}^{l}_{n}\mid\mathbf{b}^{l}_{b}(\mathbf{x}_{n})\boldsymbol{\mu}_{b}^{l},\tilde{\sigma}^{2}_{n}\right)-\frac{1}{2\tilde{\sigma}^{2}_{n}}\left[\tilde{k}_{nn}+\operatorname{tr}(\mathbf{A}^{l}_{b}\boldsymbol{\Lambda}_{n})\right]\right\}\\ &\quad-\text{KL}[q(\mathbf{z}^{l}_{\mathbf{U}})\mid\mid p_{\psi}(\mathbf{z}^{l}_{\mathbf{U}})],\end{aligned}$
where $\mathbf{b}^{l}_{b}(\mathbf{x})=\mathbf{K}^{l}_{\mathbf{xU}}\left(\mathbf{K}^{l}_{\mathbf{UU}}\right)^{-1}$, $\tilde{k}_{nn}=\mathbf{K}_{\mathbf{x}_{n}\mathbf{x}_{n}}-\mathbf{b}^{l}_{b}(\mathbf{x}_{n})\mathbf{K}_{\mathbf{U}\mathbf{x}_{n}}$, $\boldsymbol{\Lambda}_{n}={\mathbf{b}^{l}_{b}}^{\top}(\mathbf{x}_{n})\cdot\mathbf{b}^{l}_{b}(\mathbf{x}_{n})$, and $q(\mathbf{Z_{U}})=\prod_{l=1}^{L}\mathcal{N}(\mathbf{z}^{l}_{\mathbf{U}}\mid\boldsymbol{\mu}^{l}_{b},\mathbf{A}^{l}_{b})$. The time complexity is $\mathcal{O}(N_{b}M^{2}+M^{3})$ with $M$ inducing points.
@@
  • Compute $\mathbf{\Sigma}_{N}^{l}\in\mathbb{R}^{M\times M}$, $\mathbf{\mu}_{N}^{l}\in\mathbb{R}^{M}$ and $\mathbf{A}_{N}^{l}\in\mathbb{R}^{M\times M}$ using the estimators (18) for $l\in\{1,2,...,L\}$.
@@ D.2.1 SDE Representations for Markovian GPs
[…]
$\begin{gathered}\mathbf{F}=\begin{bmatrix}0&1\\ -\lambda^{2}&-2\lambda\end{bmatrix},\quad\mathbf{L}=\begin{bmatrix}0\\ 1\end{bmatrix},\quad\mathbf{H}=\begin{bmatrix}1&0\end{bmatrix},\\ \mathbf{m}_{0}=\mathbf{0},\quad\mathbf{P}_{0}=\begin{bmatrix}\sigma^{2}&0\\ 0&\sigma^{2}\lambda^{2}\end{bmatrix},\quad\mathbf{Q}_{c}=4\lambda^{3}\sigma^{2}\end{gathered}$ (21)
with $d=2$ and $\lambda=\frac{\sqrt{3}}{l}$, $l$ denotes the lengthscale. Furthermore, we have the following expression for $\mathbf{A}_{i,i+1}$:
$\mathbf{A}_{i,i+1}=e^{\Delta_{i}\mathbf{F}}=e^{-\Delta_{i}\lambda}\cdot\left\{\Delta_{i}\left[\begin{array}[]{cc}\lambda&1\\ -\lambda^{2}&-\lambda\end{array}\right]+\mathbf{I}\right\}.$
[…]
Filtering and Smoothing in Linear Gaussian SSMs
Consider a linear Gaussian SSM with system states $\{\mathbf{s}_{t}\in\mathbb{R}^{d}\}_{t=1}^{T}$ and observations $\{\mathbf{y}_{t}\in\mathbb{R}^{K}\}_{t=1}^{T}$  [^3] Some notations in this subsection may be inconsistent to the previous part. The goal is to show the algorithm as neatly as possible.:
$\mathrm{initial\ distribution:\quad}$ $p(\mathbf{s}_{0})=\mathcal{N}(\mathbf{s}_{0}\mid\mathbf{m}_{0},\mathbf{P}_{0})$
$\mathrm{Markovian\ dynamics:\quad}$ $p(\mathbf{s}_{t}\mid\mathbf{s}_{t-1})=\mathcal{N}(\mathbf{s}_{t}\mid\mathbf{A}_{t-1}\mathbf{s}_{t-1}+\mathbf{b}_{t-1},\mathbf{Q}_{t-1})$
$\mathrm{(linear\ Gaussian)\ measurements:\quad}$ $p(\mathbf{y}_{t}\mid\mathbf{s}_{t})=\mathcal{N}(\mathbf{y}_{t}\mid\mathbf{H}_{t}\mathbf{s}_{t}+\mathbf{c}_{t},\mathbf{R}_{t})$
The quantities of interest are the prediction $p(\mathbf{s}_{t}\mid\mathbf{y}_{1:t-1})$, the filtering $p(\mathbf{s}_{t}\mid\mathbf{y}_{1:t})$, the data likelihood $p(\mathbf{y}_{t}\mid\mathbf{y}_{1:t-1})$, and the smoothing $p(\mathbf{s}_{t}\mid\mathbf{y}_{1:T})$. All of these terms have closed forms in the linear Gaussian setting.
$\mathrm{prediction:\quad}$ $p(\mathbf{s}_{t}\mid\mathbf{y}_{1:t-1})=\mathcal{N}(\mathbf{s}_{t}\mid\mathbf{m}^{p}_{t},\mathbf{P}^{p}_{t})$
$\mathrm{filtering:\quad}$ $p(\mathbf{s}_{t}\mid\mathbf{y}_{1:t})=\mathcal{N}(\mathbf{s}_{t}\mid\mathbf{m}^{f}_{t},\mathbf{P}^{f}_{t})$
$\mathrm{data\ likelihood:\quad}$ $p(\mathbf{y}_{t}\mid\mathbf{y}_{1:t-1})=\mathcal{N}(\mathbf{y}_{t}\mid\hat{\mathbf{y}}_{t},\mathbf{S}_{t})$
$\mathrm{smoothing:\quad}$ $p(\mathbf{s}_{t}\mid\mathbf{y}_{1:T})=\mathcal{N}(\mathbf{s}_{t}\mid\mathbf{m}^{s}_{t},\mathbf{P}^{s}_{t})$
The SSM inference includes a forward (i.e., Kalman filter) and a backward (i.e., smoothing) process. The updated equations are given by
  • Kalman filter (with $\mathbf{m}^{f}_{0}=\mathbf{0},\mathbf{P}^{f}_{0}=\mathbf{P}_{0}$)
    $p(\mathbf{s}_{t}\mid\mathbf{y}_{1:t-1}):$$\begin{cases}\mathbf{m}^{p}_{t}=\mathbf{A}_{t-1}\mathbf{m}^{f}_{t-1}+\mathbf{b}_{t-1},\\ \mathbf{P}^{p}_{t}=\mathbf{A}_{t-1}\mathbf{P}^{f}_{t-1}\mathbf{A}^{\top}_{t-1}+\mathbf{Q}_{t-1}.\end{cases}$
    $p(\mathbf{y}_{t}\mid\mathbf{y}_{1:t-1}):$$\begin{cases}\hat{\mathbf{y}}_{t}=\mathbf{H}_{t}\mathbf{m}^{p}_{t}+\mathbf{c}_{t},\\ \mathbf{S}_{t}=\mathbf{H}_{t}\mathbf{P}^{p}_{t}\mathbf{H}_{t}^{\top}+\mathbf{R}_{t}.\end{cases}$
    $p(\mathbf{s}_{t}\mid\mathbf{y}_{1:t}):$$\begin{cases}\mathbf{m}^{f}_{t}=\mathbf{m}^{p}_{t}+\mathbf{W}_{t}(\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}),\\ \mathbf{P}^{f}_{t}=\mathbf{P}^{p}_{t}-\mathbf{W}_{t}\mathbf{S}_{t}\mathbf{W}^{\top}_{t},\end{cases}\mathrm{where\ }\mathbf{W}_{t}=\mathbf{P}^{p}_{t}\mathbf{H}_{t}^{\top}\mathbf{S}^{-1}_{t}.$
  • Rauch-Tung-Striebel(RTS) smoother (with $\mathbf{m}^{s}_{T}=\mathbf{m}^{f}_{T}$)
    $p(\mathbf{s}_{t}\mid\mathbf{y}_{1:T}):\begin{cases}\mathbf{m}^{s}_{t}=\mathbf{m}^{f}_{t}+\mathbf{G}_{t}(\mathbf{m}^{s}_{t+1}-\mathbf{m}^{p}_{t+1}),\\ \mathbf{P}^{s}_{t}=\mathbf{P}^{f}_{t}+\mathbf{G}_{t}(\mathbf{P}^{s}_{t+1}-\mathbf{P}^{p}_{t+1})\mathbf{G}_{t}^{\top},\end{cases}\mathrm{where}\ \mathbf{G}_{t}=\mathbf{P}^{f}_{t}\mathbf{A}^{\top}_{t}\left[\mathbf{P}^{p}_{t+1}\right]^{-1}.$
The above Bayesian smoothing algorithm has linear complexity regarding the sequence length $T$.

### v3

@@ 1 Introduction
In this work, we propose an efficient neighbour-driven approach to GPVAE training and prediction (as illustrated in Fig. 1). Our key insight is that in many structured datasets—such as video frames with temporal adjacency or spatial regions with local patterns—focusing on a small set of nearest neighbours captures most of the essential correlation structures. This intuition aligns with the “first law of geography,” which underlies various Nearest Neighbour GP (NNGP) methods (Vecchia, 1988; Stein et al., 2004). Inspired by recent evidence that NNGPs can outperform sparse GPs with inducing points (Hensman et al., 2013) in large-scale geostatistics and machine learning tasks (Datta et al., 2016b; Tran et al., 2021b; Wu et al., 2022; Cao et al., 2023; Allison et al., 2024), we introduce two latent GP prior approximations to neighbour-driven latent modelling in GPVAEs: (1) Hierarchical Prior Approximation (HPA): enforces sparsity in dense GP covariance matrices via a hierarchical mechanism that selects local neighbourhoods in each mini-batch; (2) Sparse Precision Approximation (SPA): chain-factorises the GP prior into conditional distributions involving only the nearest neighbours. Both methods support structured latent modelling with flexible kernels while avoiding large sets of inducing points or rigid kernel assumptions. By integrating these approximations into mini-batch amortised inference, our approach leverages local correlation principles for scalable GPVAE inference, preserving essential latent structures with minimal overhead.
@@ 3.2 Neighbour-Driven Inference
where $\mathbf{D}_{\mathbf{w}}=\operatorname{diag}(\mathbf{w})$. From (6), the correlations of the unselected variables $\mathbf{Z}_{\backslash\mathcal{I}}$ among $\mathbf{Z}$ are eliminated, where $\backslash\mathcal{I}=\{i:\mathbf{w}_{i}=0\}$. Then, we set up an amortised variational distribution that uses a similar form to (6) as follows [^1] We have omitted the conditioning variable $\mathbf{Y}$.
@@
$\quad\quad\quad\quad\biggl{.}-\frac{1}{N}\mathrm{KL}\left[q(\mathbf{Z}_{n(i)})\mid\mid p(\mathbf{Z}_{n(i)})\right]\biggr{\}},$
@@
Figure 2: Latent representation learning for the moving ball dataset. (a) The leftmost column shows frames overlaid and shaded by time. The orange reconstructed paths are obtained using $M/H=10$. (b) The results of the standard VAE and two full-batch baselines are shown with the shaded bands. The error bars and the shades indicate $\pm 1$ standard deviation.
@@ 3.3 Predictive Posterior
Predictive posterior distributions $p(\mathbf{y}_{*}\mid\mathbf{x}_{*},\mathbf{Y})$ are often used in tasks of conditional generation, where an unseen auxiliary $\mathbf{x}_{*}\in\mathbb{R}^{D}$ is given. Here, we derive an approximate predictive posterior for the proposed GPVAEs based on neighbouring latent representations. Specifically, computing $p(\mathbf{y}_{*}\mid\mathbf{x}_{*},\mathbf{Y})$ at a new location $\mathbf{x}_{*}$ only needs to consider its $H$ nearest neighbours in $\mathbf{X}$ (denoted by the index set n()n(*)):
@@ 5.2 Rotated MNIST
Corrupted
Frames
VAE (Kingma & Welling, 2014) 0.193 $\pm$ 0.038 0.333 $\pm$ 0.011 75.741 $\pm$ 0.660
@@
Missing
Frames
SVGPVAE-M10 (Jazbec et al., 2021) 87.711 $\pm$ 0.484 6.122 $\pm$ 0.015 8.280 $\pm$ 0.111
@@
This task uses the sequences created by Krishnan et al. (2015). The dataset has 50,000/10,000 training/test sequences, each containing 10 frames. The rotations between two consecutive frames are normally distributed, with around 60% pixels absent randomly. Fig. 3(a) shows a sequence from the dataset.
@@ 5.4 Geostatistical Datasets
[…]
@@ Acknowledgements
We thank the reviewers for their insightful comments and constructive suggestions. We also acknowledge the CSF3 at the University of Manchester for providing GPU resources. Xinxing is supported by the UoM-CSC Joint Scholarship, and Xiaoyu receives support through the UoM Departmental Studentship for the Department of Computer Science.
@@
This paper presents work aimed at advancing machine learning. There are potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ B.1 GPVAE-HPA ELBO
We begin by outlining the derivation for GPVAE-HPA. In each training iteration, randomly sample a mini-batch $\mathcal{B}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i\in\mathcal{I}}$ with the index set $\mathcal{I}$. For each training point $(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathcal{B}$, look for the $H$ locations $\mathbf{X}_{n(i)}\in\mathbb{R}^{H\times D}$ in $\mathbf{X}=[\mathbf{x}_{n}]_{n=1}^{N}$ such that $n(i)$ are indices of top $H$ nearest points to $\mathbf{x}_{i}$. This neighbour-driven strategy materialises the sampling of a binary indicative vector $\mathbf{w}$ for latent variables. Then, the ELBO is given by
@@ B.2 GPVAE-SPA ELBO
[…]
@@ C.1 Moving Ball
Frame size $\times 32$
@@ C.2.2 Missing Frame Generation
Decoder
16 $\rightarrow$ 1568 $\rightarrow$ Deconv(32, 64, ks=3) $\rightarrow$
Deconv(64, 32, ks=3) $\rightarrow$ Deconv(32, 1, ks=3) $\rightarrow$ Sigmoid
@@ C.4.1 Jura
Table 12 presents the results on the Jura dataset discussed in the main text. We also illustrate the prediction in Fig. 9.
@@ Appendix D Scalable Gaussian Process Variational Autoencoders
Table 15: A summary of relevant models. NN, $M$, and $H$ are the number of data points, inducing points, and nearest neighbours. $N_{b}$ is the mini-batch size. $P$ represents the feature dimensions of low-rank representations. $d$ is the dimensions of the state variables.
Models Mini-batching* GP complexity
Arbitrary
kernel
Standard VAE
encoder
Reference
@@ D.1 SVGPVAE
Replacing exact GPs with Sparse Variational GP (SVGP) techniques (Titsias, 2009; Hensman et al., 2013) to enhance the scalability of GPVAEs is a natural approach. However, such methods cannot be directly integrated into GPVAEs, as they either require the entire dataset to be loaded into memory or are not suited for amortization. In response to this scalability challenge, Jazbec et al. (2021) propose a novel approach that leverages a specific parameterisation of the distribution of inducing variables based on Pearce (2020). This formulation enables the training objective to accommodate mini-batching, amortization, and the use of arbitrary kernels.
@@
where $Z^{l}_{\psi,\phi}=\mathcal{N}(\mu^{l}_{\phi}(\mathbf{Y})\mid\mathbf{0},k^{l}_{\psi}(\mathbf{X},\mathbf{X})+{\sigma^{2}_{\phi}}^{l}(\mathbf{Y}))$. The posterior $q_{\psi,\phi}(\mathbf{z}_{n}\mid\mathbf{y}_{n})$ can be regarded as derived from a “pseudo” GP regression with input-observation pairs $(\mathbf{X},\mu^{l}_{\phi}(\mathbf{Y}))$ and noise ${\sigma_{\phi}^{2}}^{l}(\mathbf{Y})$. We denote the latent dataset as $\{\mathbf{X},\tilde{\mathbf{Y}}:=\mu_{\phi}(\mathbf{Y}),\tilde{\boldsymbol{\sigma}}^{2}:={\sigma_{\phi}^{2}}^{l}(\mathbf{Y})\}$. This training objective will require $\mathcal{O}(N^{3})$ time complexity due to $\log Z^{l}_{\psi,\phi}(\mathbf{Y},\mathbf{X})$ and $q_{\psi,\phi}(\mathbf{z}_{n}\mid\mathbf{y}_{n})$ terms without any approximation. Jazbec et al. (2021) approximate the encoder’s output distribution $q_{\psi,\phi}(\mathbf{z})$ using SVGP’s posterior:
@@
$\mathbf{\Sigma}_{b}^{l}$ $=\mathbf{K_{UU}}+\frac{N}{N_{b}}\mathbf{K}_{\mathbf{UX}_{b}}\operatorname{diag}(\tilde{\boldsymbol{\sigma}}_{b}^{-2})\mathbf{K}_{\mathbf{X}_{b}\mathbf{U}},$ (18)
$\boldsymbol{\mu}_{b}^{l}$ $=\frac{N}{N_{b}}\mathbf{K_{UU}}(\mathbf{\Sigma}_{b}^{l})^{-1}\mathbf{K}_{\mathbf{U}\mathbf{X}_{b}}\operatorname{diag}(\tilde{\boldsymbol{\sigma}}_{b}^{-2})\tilde{\mathbf{Y}}^{l}_{b},$
$\mathbf{A}_{b}^{l}$ $=\mathbf{K_{UU}}(\mathbf{\Sigma}_{b}^{l})^{-1}\mathbf{K_{UU}}.$
@@
  • Compute $\mathbf{\Sigma}_{N}^{l}\in\mathbb{R}^{M\times M}$, $\boldsymbol{\mu}_{N}^{l}\in\mathbb{R}^{M}$ and $\mathbf{A}_{N}^{l}\in\mathbb{R}^{M\times M}$ using the estimators (18) for $l\in\{1,2,...,L\}$.
@@ D.2.1 SDE Representations for Markovian GPs
Note that $d,\mathbf{F},\mathbf{L},\mathbf{H},\mathbf{m}_{0},\mathbf{P}_{0},\mathbf{Q}_{c}$ in (19) and (20) depends on the kernel’s type and parameters.